# AMEX: Android Multi-annotation Expo Dataset for Mobile GUI Agents

### v1

@@ A.1.3 GUI screen and element functionality description collection details
The prompt is in the following format.
@@ 4. Output Format:
A.1.4 GUI-Action chain collection details
@@
A.1.5 Collection resources details
@@
  • GUI screen and element descriptions use GPT-4o API, which consumes about 600 dollars.
  • Instructions with GUI-action chains take approximately 200 human-hour.
@@ A.2 Experiment details
In our implementation, we utilize the internlm-7b variant of the SPHINX-X model, as detailed in [7]. The pre-trained checkpoint for this model was sourced from the official repository mentioned in [21]. For image processing, the input images, each sized $1024\times 1024$, are segmented into sub-images. Visual features from these sub-images are extracted using two distinct visual encoders: DINOv2 [18] and ConvNext [25]. To ensure compatibility in feature dimensions across different modules, linear projection layers are employed to align the channel dimensions. Regarding the model’s parameter settings, as outlined in Section 4, we configure the history window size to four. Additionally, we introduce a special token, <ICON>, specifically designed to identify interactive elements within the interface, strengthening the model’s interpretability and responsiveness to user interactions. The agent is trained on a cluster with 3 nodes, each with eight NVIDIA A100 (80GB) GPUs. The fine-tuning was completed in four epochs.
@@ A.3 More AMEX examples
See Figure 9 for more examples of GUI interactive elements grounding and description.
See Figure 10 and Figure 11 for more examples of instruction with GUI-action chains.

### v2

@@ A.1.4 GUI screen and element functionality description collection details
The prompt is in the following Listing 1 format.
@@ 4. Output Format:
A.1.5 GUI-Action chain collection details
The instruction generation process comprises three phases. Initially, human annotators create 5-10 complex instructions for each target app, considering its specific functions and capabilities. These initial instructions, combined with relevant app metadata collected online, serve as input for GPT. It then generates a larger set of 80-100 instructions that exhibit similar structure and intent to the human-provided examples. However, due to GPT’s limitations on understanding real-world constraints, human filtering is adopted to modify or remove any unreasonable or impractical instructions. For example, the human-proposed instruction “Open Booking.com, search for stays in Washington DC from 2024 June 11 to June 15.” would generate a similar but impractical “Open Booking.com, search for stays in New York from 2024 May 3 to May 7”, which needs modification since the dates are not selectable when the task is generated.
Since the meta information online doesn’t reflect all functions of an app and sometimes may include misleading information, the instruction generated by GPT may have two types of impracticabilities: (i) the task is unrelated to the app activity (ii) some of the steps in the instruction are impractical. For instance, an app “SEPHORA” only sells beauty related products, but GPT may generate tasks that searches for electronics. Another example would be about a ticket booking app “SeekGeek”, which doesn’t provide a sort option “sort by price from high to low”. While human created task would include “sort by price from low to high”, GPT would generate something with“sort by price from high to low”. This step is impractical so we would filter it out.
@@ A.1.5 GUI-Action chain collection details
A.1.6 Collection resources details
@@
  • GUI screen and element descriptions use GPT-4o and Gemini API, which consumes about 800 dollars.
  • Instructions with GUI-action chains take approximately 300 human-hour.
@@ A.2 Experiment details
In our implementation, we utilize the internlm-7b variant of the SPHINX model, as detailed in Liu et al. (2024). The pre-trained checkpoint for this model was sourced from the official repository mentioned in Team (2023). For image processing, the input images, each sized $1024\times 1024$, are segmented into sub-images. Visual features from these sub-images are extracted using two distinct visual encoders: DINOv2 Oquab et al. (2024) and ConvNext Woo et al. (2023). To ensure compatibility in feature dimensions across different modules, linear projection layers are employed to align the channel dimensions. Regarding the model’s parameter settings, as outlined in Section 4.1, we configure the history window size to four. Additionally, we introduce a special token, <ICON>, specifically designed to identify interactive elements within the interface, strengthening the model’s interpretability and responsiveness to user interactions. The agent is trained on a cluster with 3 nodes, each with eight NVIDIA A100 (80GB) GPUs. The fine-tuning was completed in four epochs.
A.2.1 AndroidControl details
AndroidControl splits the test set into four categories, i.e., in-domain data (IDD), category unseen, app unseen, and task unseen. Besides, the instructions are split into two levels. The high-level instructions are simply the goal of the task and the low-level instructions are the guides for each step during the task execution. Low-level tasks are easier for agents due to the fact that the agent only need to perform the action specified by the step-wise low-level instruction, while the high-level tasks require the agents to perform actions based on the overall instruction. For example, a step-wise low-level instruction is “click the profile element at the bottom right of the screen.”, compared with high-level instruction “Change my profile name to Unknown111”.
A.2.2 AitW details
MMR SphAgent vs. Human 90.91 83.65 87.50 87.68 82.39 86.43
AitW Anno. vs. Human 88.18 75.00 82.69 84.78 75.57 81.25
Random SphAgent vs. Human 92.31 93.75 94.33 94.83 94.55 93.95
AitW Anno. vs. Human 94.23 90.34 93.62 93.97 95.00 93.43
We observe that there are several types of error cases in the AitW test set (see Figure 6). To assess the unreliability of the original AitW test set annotations, we hire human annotators to evaluate two subsets derived from the original test set. One subset is randomly chosen from episodes where SphAgent receives low accuracy (named Mis-Match Random), i.e., about 66% compared to the overall 78.72% shown in Table 5, indicating a high mismatch between inference results and AitW annotations. The other subset is randomly selected from the remaining episodes (named Random). Annotators first filter out repeated and redundant screenshots (see Figure 6(a)). For each remaining screenshot, the annotators then record the accuracies of SphAgent-inferred actions and the original AitW annotations versus our human evaluations. Figure 7 illustrates cases where annotators mark both the inferred action and the original annotation as correct, even though they interact with different elements. Table 7 presents the accuracy from human evaluation. The table indicates that the low SphAgent accuracy in the MMR subset are primarily due to unsatisfactory annotations (AitW Anno), which are used as ground truth during the evaluation. Both SphAgent and AitW Anno in the “MMR” subset have lower accuracies than those in the “Random” subset, highlighting that the tasks in the MMR subset are relatively more challenging. Furthermore, SphAgent achieves better human evaluation results than the original annotations, demonstrating its effectiveness and close alignment with human judgment. Comparing the results from Table 5 and Table 7, the notable differences in overall scores (i.e., 78.72%, 86.43%, and 93.95%) underscore the unreliability and misleading nature of the AitW test set annotations.
(a) Repeating useless screenshots and actions.
(b) Wrong task complete.
@@ A.3 More AMEX examples
See Figure 8 for more examples of GUI interactive elements grounding and description. See Figure 9 for more examples of instruction with GUI-action chains.